OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (7, -2) and has a slope of -3. y - 2 = -3(x + 7) y + 7 = -3(x - 2) y + 2 = -3(x - 7) y - 7 = -3(x + 2)

OpenStudy (mathstudent55):

The point-slope equation of a line is: \( y - y_1 = m(x - x_1) \) where m = slope and the point is \( (x_1, y_1) \)

OpenStudy (mathstudent55):

Place your given slope and point in the equation above and see which equation it is equal to.

OpenStudy (anonymous):

So its b

OpenStudy (mathstudent55):

correct